given a typical household uses 20kwh of energy in a single day,what charge capacity would be an energy storage system need to give it would run at 2.5 V redox flow battery?
Answer #1

Given EnergyE=20Kwh

we know Charge Q=i*t

E=V*I*t

20*103 =2.5 *i*t

2.22=Q

Q=2.22 Columbs